### v1

@@ 7 Discussion
As far as non-abelian NS5-brane dynamics is concerned, the addition of fundamental string and momentum charges is just window dressing, and has no effect on W-brane energetics – it is as if one were simply working with fivebranes on the Coulomb branch. As a result, in the fivebrane decoupling limit (1.2) the non-abelian scale is reached when the scaled fivebrane separation is of order $\ell_{na}$, Eq. (5.3) (i.e. order one if the ${\mathbb{T}}^{4}$ is string scale in size), so that the tension of little strings (2.51) is order one.
@@ 7.1 The quantum scale and the non-abelian scale
The near-horizon region of a nearly-BPS black hole is subject to large quantum fluctuations. A universal sector of the excitation spectrum of such black holes is the set of Schwarzian modes of the nearly-$AdS_{2}$ throat (see for instance Ghosh:2019rcj ; Heydeman:2020hhw ; Iliesiu:2020qvm ; Lin:2022rzw ; Kolanowski:2024zrq ). As a BTZ black hole approaches extremality, these modes undergo large quantum fluctuations. The Schwarzian excitations are concentrated very near the horizon (see e.g. the explicit solution in Kolanowski:2024zrq ), and are thus parametrically lighter than other modes in the throat; they largely decouple at low energies from other modes, and can be regarded as fluctuations in the length of the throat.
The large quantum fluctuations in the depth of the throat at extremality lead to a finite effective depth of the throat, and an effective “stretched horizon” at a radius we have called the quantum scale $r_{\rm qu}$. Quantum fluctuations of fivebrane stars become large when fivebrane strands approach within the non-abelian scale $\ell_{na}=\mu\big(\frac{V_{4}}{{\alpha^{\prime}}^{2}}\big)^{\!\frac{1}{2}}$. Since the onset of black hole formation is associated to a deconfinement transition in the underlying brane dynamics, it is natural to identify these two scales, and indeed we have seen in examples that the $AdS_{2}$ throat lengths associated to these scales match.
@@
${\mathsf{F}}_{\mathsf{m}}=e^{-\lambda}{{\mathbb{F}}}_{m}~~,~~~~{\mathsf{X}}_{\mathsf{n}}=e^{-\lambda}{\mathbb{X}}_{\mathsf{n}}~.$ (7.1)
@@ 7.3 Near-BPS chaos
The Thouless time is the inverse of the Thouless energy, which characterizes the extent of the repulsion between energy levels. The much larger entropy of three-charge fivebrane stars relative to their two-charge counterparts, and a much more chaotic near-BPS dynamics, should result in a much shorter Thouless time as compared to the two-charge case. The transition to maximal chaos should be associated to horizon formation, which happens when the system begins to access its fully nonabelian excitations – in the present context, the little strings. The rapid reconnection of the little strings results in fast scrambling, which again is the hallmark of the system being able to rapidly access distant regions of the available configuration space Sekino:2008he . The weaker chaos of the fivebrane star dynamics is associated to the longer time a chaotically diffusing string or graviton takes to explore its environment, interact with its neighbors and make an appreciable change in the state of the system (for instance, to alter the fivebrane supertube configuration). It would be interesting to explore this chaotic dynamics and determine its characteristics
@@ 7.5 Near-BPS gravitational collapse
As discussed in Martinec:2019wzw ; Martinec:2024emf , at some point these new, stringy degrees of freedom become light enough that they can no longer be ignored. These are the “little strings” of non-abelian fivebrane dynamics, which have been suppressed by self-consistently holding the fivebranes apart via their transverse excitation. Turning off those excitations, or transferring them to other parts of the system, allows the fivebranes to come together and liberate the Hagedorn phase of little strings that accounts for black fivebrane entropy. In this way, we see in the bulk description the relation between horizon formation (the Hawking-Page transition) and the deconfinement transition of the non-abelian brane dynamics Witten:1998zw .
@@ C.2 BPS ground states
$n_{y}=-\bigg(\frac{h_{L}+M^{\prime}w^{\prime}+\frac{{n_{5}}}{4}w^{\prime 2}}{w_{y}}\bigg)~~,~~~~M=-M^{\prime}-\frac{{n_{5}}}{2}w^{\prime}+{\mathsf{p}}\,w_{y}~.$ (C.9)
Note that the T-dual of superstratum excitations has $h_{L}=w^{\prime}=0$ in order to be straight strings carrying no momentum. This makes sense, because SU(2){SU(2)} spectral flow is an internal automorphism of the representations, and thus can be re-expressed in terms of current raising operators (and thus oscillator excitations).
@@
$\mathcal{P}_{u}$ $=0~~,~~~~\mathcal{P}_{v}=-n_{y}+\frac{2j{\mathsf{p}}}{{n_{5}}}~~,~~~~\mathcal{J}_{R}=0~~,~~~~\mathcal{J}_{L}=\big(M^{\prime}+j\big)+\frac{{n_{5}}}{2}w^{\prime}$ (C.12)
where we have used (C.9). Vanishing Killing momentum $\mathcal{P}_{u}$ is a requirement for BPS-ness; $\mathcal{J}_{R}=0$ agrees with the absence of a string contribution to right-moving angular momentum in (6.1). Below, we will match $\mathcal{P}_{v}$ to (6.1) using (C.9), after determining the string ring size $\mathfrak{b}$} in terms of the worldsheet quantum numbers. Note the additional contribution $\frac{2j{\mathsf{p}}}{{n_{5}}}$ as compared to the naively expected first term. As discussed in Martinec:2025xoy , the gauge invariant charge (C.12) compares the string charge (here momentum charge, there winding charge) to the corresponding charge carried by the fivebrane for a given charge budget. This subtraction is in general not quantized, but renders the expression for the charge invariant under large gauge tranformations (C.8), which trade charge carried on the string for charge carried as flux in the fivebrane background. This additional term should be dropped when comparing to the corresponding spacetime quantity.
@@
$\Big(\frac{r_{1}}{{\mathfrak{a}}}\Big)^{2}=\frac{M-j}{2j}~~,~~~~\sin^{2}\theta_{1}=\frac{j-M^{\prime}}{2j}$ (C.13)
up to shifts of order $1/j$ stemming from the effects of the string polarization state, with the width of the peak scaling as $j^{-1/2}$.
@@ C.3 Classical strings
$\cosh\rho_{1}=\cosh(\alpha/2)~~\Longrightarrow~~r_{1}=a\sinh\big({\textstyle\frac{1}{2}}\alpha\big)~.$ (C.18)
@@
$g_{\text{su}}=\left(\begin{matrix}e^{-i\phi}\sin\theta~&~e^{i\psi}\cos\theta\\[2.84544pt] -e^{-i\psi}\cos\theta~&~e^{i\phi}\sin\theta\end{matrix}\right)=\left(\begin{matrix}e^{-i[(2\nu^{\prime}-w^{\prime})\xi_{0}-w^{\prime}\xi_{1}]/2}\cos\frac{\alpha^{\prime}}{2}~&~e^{+i[(2\nu^{\prime}+w^{\prime})\xi_{0}+w^{\prime}\xi_{1}]/2}\sin\frac{\alpha^{\prime}}{2}\\[2.84544pt] -e^{-i[(2\nu^{\prime}+w^{\prime})\xi_{0}+w^{\prime}\xi_{1}]/2}\sin\frac{\alpha^{\prime}}{2}~&~e^{+i[(2\nu^{\prime}-w^{\prime})\xi_{0}-w^{\prime}\xi_{1}]/2}\cos\frac{\alpha^{\prime}}{2}\end{matrix}\right)~,$ (C.19)
@@
$\sin\theta_{1}=\cos(\alpha^{\prime}/2)~~\Longrightarrow~~\theta_{1}=\frac{1}{2}\big(\pi-\alpha^{\prime}\big)~.$ (C.20)
@@
${\mathsf{X}}^{1}+i{\mathsf{X}}^{2}$ $=\sqrt{r_{1}^{2}+{\mathfrak{a}}^{2}}\,\sin\theta_{1}\,{\rm exp}\Big[-\frac{i\sqrt{2}w^{\prime}v}{2w_{y}R_{y}}\Big]~~,~~~~{\mathsf{X}}^{3}+i{\mathsf{X}}^{4}=r_{1}\,\cos\theta_{1}\,{\rm exp}\Big[+\frac{i\sqrt{2}w^{\prime}v}{2w_{y}R_{y}}\Big]~.$ (C.24)
@@ C.4 Comparing to superstrata
$k=-\frac{w^{\prime}}{2}~.$ (C.25)
Also, in the comparison to the effective single-mode superstrata of section 6.2, one has
$\ell=2j~~,~~~~m+n=w_{y}~~,~~~~\kappa={\mathsf{p}}~.$ (C.26)
Two rather useful relations for strings lying in the ${\mathsf{x}}^{1}$- ${\mathsf{x}}^{2}$ plane are
@@
Inserting them in the expression for the smeared source angular momentum $J_{L}$, (6.1), the string contribution to the angular momentum charge $J_{L}$ in (6.1) evaluates to
$J_{L,1}=\frac{n_{1}}{w_{y}}\,\frac{{\mathsf{p}}w_{y}(j+M^{\prime})+kn_{5}(M-j)}{M+M^{\prime}}=\frac{n_{1}}{w_{y}}\,\frac{{\mathsf{p}}w_{y}m+kn_{5}n}{w_{y}}~.$ (C.28)
For strings lying in the ${\mathsf{x}}^{1}$- ${\mathsf{x}}^{2}$ plane, either ${\mathfrak{b}}>{\mathfrak{a}}$, in which case $m=0$; or ${\mathfrak{b}}<{\mathfrak{a}}$, in which case $n=0$. While this expression was derived using strings restricted to this plane, when expressed in terms of the quantum numbers of the general string solutions of section C.2, it holds for strings in general position. We find agreement between the angular momentum $J_{L}$ in the smeared supergravity solution (6.1) and from the exactly solvable worldsheet dynamics (C.12). These examples involve spiraling string sources that carry momentum as well as winding along ${\mathbb{S}}^{1}_{y}$; when specialized to strings having winding but carrying no momentum, we find agreement with the angular momentum of the supergravity wave of single-mode superstrata.
The relations (C.27) also allow us to write the momentum charge (6.1) as
$n_{p}$ $=\bigg(1+\frac{{n_{1}}\mu^{2}}{|{\mathfrak{b}}^{2}-{\mathfrak{a}}^{2}|}\bigg)\frac{{\mathsf{p}}^{2}{\mathfrak{a}}^{2}}{n_{5}\mu^{2}}+\frac{{n_{1}}}{w_{y}}\bigg[\frac{-2M^{\prime}k+{n_{5}}k^{2}}{w_{y}}+\frac{k{\mathsf{p}}(j+M^{\prime})}{kn_{5}+{\mathsf{p}}w_{y}}\bigg]$
$=\frac{{\mathsf{p}}\,2J_{R}}{{n_{5}}}+\frac{{n_{1}}}{w_{y}}\bigg[\frac{{\mathsf{p}}\ell}{{n_{5}}}+\frac{{\mathsf{p}}km}{m+n}+\frac{(\ell-2m)k+{n_{5}}k^{2}}{w_{y}}\bigg]$ (C.29)
@@
To compare to superstrata, we set $k=0$, then the first term on the R.H.S. of the first line reproduces (6.44). We find the T-dual of the “strand budget” relation (6.38) of superstrata
@@
Momentum quantization on individual strands imposes $n_{1}\in w_{y}{\mathbb{Z}}$. In the original superstratum frame, the third charge is momentum, $n_{1}=\tilde{n}_{\tilde{p}}=(m+n)N_{0}$ where $N_{0}$ is the number of internal R-R scalar modes excited on the fivebrane, Eq. (6.46); the background strands have length $\mathsf{p}$}, and so their contribution to the string winding budget in that frame is ${\mathsf{p}}\ell N_{0}$, accounting for the second term on the R.H.S. The first term on the R.H.S. is the contribution to the strand budget of the underlying fivebrane supertube, from modes which T-dualize to the part of the string winding budget coming from the number $N_{+}=2J_{R}$ of transversely polarized excitations on the fivebrane, each of strand length $\mathsf{p}$} and carrying a half unit of angular momentum on both left and right.
We thus have a complete match between (1) the string profiles ${\mathsf{X}}(v)$ in our classical solutions for round sources elaborated in section 6.1; and (2) BPS ground states in the exact worldsheet theory of the round NS5-P supertube, in which the strings carry no momentum. We also have a generalization to backgrounds generated by stringy soures, carrying both winding and momentum, when $k\neq 0$.
@@
The argument of the square root is $n_{1}n_{5}n_{p}$, while the remaining factor in the argument of the log is smaller than one – affine SU(2){SU(2)} highest weights are restricted to $\ell<{n_{5}}$; and ${\mathfrak{a}}^{2}=\frac{2J_{R}}{{\mathsf{p}}}\mu^{2}$, so for the fivebrane ring to be larger than the quantum scale one must have $2J_{R}>{\mathsf{p}}$. Thus, once again the length of the throat is at most that of the extremal black hole, Eq. (2.43).

### v2

@@ 7 Discussion
As far as non-abelian NS5-brane dynamics is concerned, the addition of fundamental string and momentum charges is just window dressing, and has no effect on W-brane energetics – it is as if one were simply working with fivebranes on the Coulomb branch. As a result, in the fivebrane decoupling limit (1.2) the non-abelian scale is reached when the scaled fivebrane separation is of order $\hat{\ell}_{na}$, Eq. (5.3) (i.e. order one if the ${\mathbb{T}}^{4}$ is string scale in size), so that the tension of little strings (2.51) is order one.
@@ 7.1 The quantum scale and the non-abelian scale
The near-horizon region of a nearly-BPS black hole is subject to large quantum fluctuations. A universal sector of the excitation spectrum of such black holes is the set of Schwarzian modes of the nearly-$AdS_{2}$ throat (see for instance Ghosh:2019rcj ; Heydeman:2020hhw ; Iliesiu:2020qvm ; Lin:2022rzw ; Kolanowski:2024zrq ; Castro:2025itb ; Acito:2025hka ). As a BTZ black hole approaches extremality, these modes undergo large quantum fluctuations. The Schwarzian excitations are concentrated very near the horizon (see e.g. the explicit solution in Kolanowski:2024zrq ), and are thus parametrically lighter than other modes in the throat; they largely decouple at low energies from other modes, and can be regarded as fluctuations in the length of the throat.
The large quantum fluctuations in the depth of the throat at extremality lead to a finite effective depth of the throat, and an effective “stretched horizon” at a radius we have called the quantum scale $r_{\rm qu}$. Quantum fluctuations of fivebrane stars become large when fivebrane strands approach within the non-abelian scale $\hat{\ell}_{na}^{2}=\big(\frac{V_{4}}{{\alpha^{\prime}}^{2}}\big)^{\!\frac{1}{2}}$. Since the onset of black hole formation is associated to a deconfinement transition in the underlying brane dynamics, it is natural to identify these two scales, and indeed we have seen in examples that the $AdS_{2}$ throat lengths associated to these scales match.
@@
${\mathsf{F}}_{\mathsf{m}}=e^{-\lambda}{\mathbb{F}}_{\mathsf{m}}~~,~~~~{\mathsf{X}}_{\mathsf{n}}=e^{-\lambda}{\mathbb{X}}_{\mathsf{n}}~.$ (7.1)
@@ 7.3 Near-BPS chaos
The much larger entropy of three-charge fivebrane stars relative to their two-charge counterparts, and a much more chaotic near-BPS dynamics, should result in a much shorter Thouless time as compared to the two-charge case. The transition to maximal chaos should be associated to horizon formation, which happens when the system begins to access its fully nonabelian excitations – in the present context, the little strings. The rapid reconnection of the little strings results in fast scrambling, which again is the hallmark of the system being able to rapidly access distant regions of the available configuration space Sekino:2008he . The weaker chaos of the fivebrane star dynamics is associated to the longer time a chaotically diffusing string or graviton takes to explore its environment, interact with its neighbors and make an appreciable change in the state of the system (for instance, to alter the fivebrane supertube configuration). It would be interesting to explore this chaotic dynamics and determine its characteristics
@@ 7.5 Near-BPS gravitational collapse
As discussed in Martinec:2019wzw ; Martinec:2024emf , at some point these new, stringy degrees of freedom become light enough that they can no longer be ignored. These are the “little strings” of non-abelian fivebrane dynamics, which have been suppressed by self-consistently holding the fivebranes apart via their transverse excitation. Turning off the momentum excitations that keep the fivebranes apart, or transferring them to other parts of the system, allows the fivebranes to come together and liberate the Hagedorn phase of little strings that accounts for black fivebrane entropy. In this way, we see in the bulk description the relation between horizon formation (the Hawking-Page transition) and the deconfinement transition of the non-abelian brane dynamics Witten:1998zw .
@@ C.2 BPS ground states
$n_{y}=-\bigg(\frac{h_{L}+M^{\prime}w^{\prime}+\frac{{n_{5}}}{4}w^{\prime 2}}{w_{y}}\bigg)~~,~~~~M=-M^{\prime}-\frac{{n_{5}}}{2}w^{\prime}+{\mathsf{p}}\,w_{y}~;$ (C.9)
momentum quantization imposes $n_{y}\in{\mathbb{Z}}$. Note that the T-dual of superstratum excitations has $h_{L}=w^{\prime}=0$ in order to be straight strings carrying no momentum. This makes sense, because SU(2){SU(2)} spectral flow is an internal automorphism of the representations, and thus can be re-expressed in terms of current raising operators (and thus oscillator excitations).
@@
$\mathcal{P}_{u}$ $=0~~,~~~~\mathcal{P}_{v}=-n_{y}+\frac{2j{\mathsf{p}}}{{n_{5}}}~~,~~~~\mathcal{J}_{R}=0~~,~~~~\mathcal{J}_{L}=\big(M^{\prime}+j\big)+\frac{{n_{5}}}{2}w^{\prime}.$ (C.12)
where we have used (C.9). Vanishing Killing momentum $\mathcal{P}_{u}$ is a requirement for BPS-ness; $\mathcal{J}_{R}=0$ agrees with the absence of a string contribution to right-moving angular momentum in (6.1). Below, we will match $\mathcal{P}_{v}$ to (6.1) using (C.9), after determining the string ring size $\mathfrak{b}$} in terms of the worldsheet quantum numbers. Note the additional fractional contribution $\frac{2j{\mathsf{p}}}{{n_{5}}}$ as compared to the naively expected integer-valued first term. As discussed in Martinec:2025xoy , the gauge invariant charge (C.12) includes a contribution to the charges carried by the string which corresponds to the amount of that charge which was carried by the fivebrane in the absence of the string, so that one maintains a fixed overall charge budget. Here this charge is momentum, and in Martinec:2025xoy it is string winding charge. This term is in general not quantized, but renders the expression for the charge invariant under large gauge transformations (C.8), which trade charge carried on the string for charge carried as flux in the fivebrane background. It arises as part of the term $\frac{|\partial_{v}{\mathsf{F}}|^{2}}{|{\mathsf{X}}-{\mathsf{F}}|^{2}}$ in the fivebrane worldvolume effective action (4). [^20] This same fractional charge contribution was noted in Giveon:2001up (see Eq. 3.36 there), and interpreted in Kim:2015gak as the worldsheet formalism for global $AdS_{3}\times{\mathbb{S}}^{3}$ working at fixed chemical potential rather than fixed charge. In the context of the gauged WZW model, which describes perturbations around the fivebrane supertube, we have a different interpretation – the worldsheet formalism conserves charge, and the fractional contributions to gauge invariant charge in correlators account for the fractional amounts of charge being carried as background flux, that are being manipulated by the vertex operator making a transition in the state of the holographic dual theory.
@@
$\Big(\frac{r_{1}}{{\mathfrak{a}}}\Big)^{2}=\frac{M-j}{2j}\equiv\frac{n}{2j}~~,~~~~\cos^{2}\theta_{1}=\frac{j+M^{\prime}}{2j}\equiv\frac{m}{2j}$ (C.13)
up to shifts of order $j^{-1}$ stemming from the effects of the string polarization state, with the width of the peak scaling as $j^{-1/2}$.
@@ C.3 Classical strings
$\cosh\rho_{1}=\cosh\big({\textstyle\frac{1}{2}}\alpha\big)~~\Longrightarrow~~r_{1}=a\sinh\big({\textstyle\frac{1}{2}}\alpha\big)~.$ (C.18)
@@
$g_{\text{su}}=\left(\begin{matrix}e^{-i\phi}\sin\theta~&~e^{i\psi}\cos\theta\\[2.84544pt] -e^{-i\psi}\cos\theta~&~e^{i\phi}\sin\theta\end{matrix}\right)=\left(\begin{matrix}e^{-i[(2\nu^{\prime}+w^{\prime})\xi_{0}+w^{\prime}\xi_{1}]/2}\cos\frac{\alpha^{\prime}}{2}~&~e^{+i[(2\nu^{\prime}-w^{\prime})\xi_{0}-w^{\prime}\xi_{1}]/2}\sin\frac{\alpha^{\prime}}{2}\\[2.84544pt] -e^{-i[(2\nu^{\prime}-w^{\prime})\xi_{0}-w^{\prime}\xi_{1}]/2}\sin\frac{\alpha^{\prime}}{2}~&~e^{+i[(2\nu^{\prime}+w^{\prime})\xi_{0}+w^{\prime}\xi_{1}]/2}\cos\frac{\alpha^{\prime}}{2}\end{matrix}\right)~,$ (C.19)
@@
$\sin\theta_{1}=\cos\big({\textstyle\frac{1}{2}}\alpha^{\prime}\big)~~\Longrightarrow~~\theta_{1}=\frac{1}{2}\big(\pi-\alpha^{\prime}\big)~.$ (C.20)
@@
${\mathsf{X}}^{1}+i{\mathsf{X}}^{2}$ $=\sqrt{r_{1}^{2}+{\mathfrak{a}}^{2}}\,\sin\theta_{1}\,{\rm exp}\Big[+\frac{i\sqrt{2}w^{\prime}v}{2w_{y}R_{y}}\Big]~~,~~~~{\mathsf{X}}^{3}+i{\mathsf{X}}^{4}=r_{1}\,\cos\theta_{1}\,{\rm exp}\Big[-\frac{i\sqrt{2}w^{\prime}v}{2w_{y}R_{y}}\Big]~.$ (C.24)
@@ C.4 Comparing to superstrata
$k=\frac{w^{\prime}}{2}$ (C.25)
so that the null constraint (C.9) reads
$m+n={\mathsf{p}}w_{y}-k{n_{5}}~.$ (C.26)
In anticipation of the relation to the effective single-mode superstrata of section 6.2, we set
$\ell=2j~.$ (C.27)
Although for the analysis here this relation can be regarded as a definition, one can verify that the string’s position (C.13) and its separation from the origin (C.14) and the fivebrane (C.15) match the corresponding features of the T-dual momentum source (6.43), with $\ell,m,n$ having identical values.
Two rather useful relations for strings lying in the ${\mathsf{x}}^{1}$- ${\mathsf{x}}^{2}$ plane follow from applying (C.16) to (C.14):
@@
Inserting them in the expression for the smeared source angular momentum $J_{L}$, (6.1), and using the null constraint (C.26) as well as (C.14), the string contribution to the angular momentum charge $J_{L}$ in (6.1) evaluates to
$J_{L,1}=\frac{n_{1}}{w_{y}}\,\frac{{\mathsf{p}}w_{y}(j+M^{\prime})+kn_{5}(M-j)}{M+M^{\prime}}=\frac{n_{1}}{w_{y}}\,\big(m+{n_{5}}k\big)~.$ (C.29)
We find agreement with the angular momentum $\mathcal{J}_{L}$ per string (C.12). Note also that there is no pathology when $m=n=0$, and the string is sitting on top of the supertube.
While the expression (6.1) was derived using strings restricted to the ${\mathsf{x}}^{1}$- ${\mathsf{x}}^{2}$ plane, when expressed in terms of the quantum numbers of the general string solutions of section C.2, it holds for strings in general position, having general $m$ and $n$. We find agreement between the angular momentum $J_{L}$ in the smeared supergravity solution (6.1) and the exactly solvable worldsheet dynamics (C.12). These examples involve spiraling string sources that carry momentum as well as winding along ${\mathbb{S}}^{1}_{y}$; when specialized to strings having winding but carrying no momentum, we find agreement with the angular momentum of the supergravity wave of single-mode superstrata.
The relations (C.14), (C.28) also allow us to write the momentum charge (6.1) as
$n_{p}$ $=\bigg(1+\frac{{n_{1}}\mu^{2}}{|{\mathfrak{b}}^{2}-{\mathfrak{a}}^{2}|}\bigg)\frac{{\mathsf{p}}^{2}{\mathfrak{a}}^{2}}{n_{5}\mu^{2}}+\frac{{n_{1}}}{w_{y}}\bigg[\frac{{n_{5}}k^{2}(\ell+n-m)}{w_{y}(m+n)}-\frac{2{\mathsf{p}}k(\ell-m)}{m+n}\bigg]$
$=\frac{{\mathsf{p}}\,2J_{R}}{{n_{5}}}+\frac{{n_{1}}}{w_{y}}\bigg[\frac{{\mathsf{p}}\ell}{{n_{5}}}+\frac{(2m-\ell)k+{n_{5}}k^{2}}{w_{y}}\bigg]$ (C.30)
@@
The term in square brackets on the second line is just $\mathcal{P}_{v}$, the momentum (C.12) of each string source. Note that again, there is no pathology when $m=n=0$, and the string sits on top of the supertube.
To compare to (orbifolded) superstrata Shigemori:2022gxf , we set $k=0$, and then the first term on the R.H.S. of the first line of (C.4) reproduces (6.45). We find the T-dual of the “strand budget” relation (6.39) of (orbifolded) superstrata
@@
There are $N_{0}=\frac{n_{1}}{w_{y}}\in{\mathbb{Z}}$ strings. In the original superstratum frame, the third charge is momentum,
$n_{1}=w_{y}N_{0}=\frac{m+n}{{\mathsf{p}}}N_{0}=\tilde{n}_{p}~,$ (C.32)
where $N_{0}$ is the number of internal R-R scalar modes excited on the fivebrane, Eq. (6.47); the background strands have length $\mathsf{p}$}, and their contribution to the string winding budget in that frame is ${\mathsf{p}}\ell N_{0}$, accounting for the second term on the R.H.S. of (C.31). The first term on the R.H.S. is the contribution to the strand budget of the underlying fivebrane supertube, from modes which T-dualize to the part of the string winding budget coming from the number $N_{+}=2J_{R}$ of transversely polarized excitations on the fivebrane, each of strand length $\mathsf{p}$} and carrying a half unit of angular momentum on both left and right.
We thus have a complete match between (1) the string profiles ${\mathsf{X}}(v)$ in our classical solutions for round sources elaborated in section 6.1; and (2) BPS ground states in the exact worldsheet theory of the round NS5-P supertube. When the strings carry no momentum, we also have a match to single-mode superstrata via T-duality, and a stringy generalization when they do carry both winding and momentum, for $k\neq 0$.
@@
The argument of the square root $n_{1}n_{5}n_{p}$ is of order the extremal entropy, while the remaining factor in the argument of the log is smaller than one – affine SU(2){SU(2)} highest weights are restricted to $\ell<{n_{5}}$; and ${\mathfrak{a}}^{2}=\frac{2J_{R}}{{\mathsf{p}}}\mu^{2}$, so for the fivebrane ring to be larger than the quantum scale one must have $2J_{R}>{\mathsf{p}}$. Thus, once again the length of the throat is at most that of the extremal black hole, Eq. (2.43).